(7) Write the equation of the line passing through (-3,1) with a slope of-1/3 Write the equation in the slope intercept form and in standard form.
Answer by checkley71(8403)

y=-1/3x+b or 1=-1/3*-3+b or 1=1+b or b=0 then y=-1/3x